1.) Suppose if the price of a good is $12, the quantity demanded is 50 units; when the price is $10, the quantity demanded is 100 units. Use the midpoint approach to compute the price elasticity of demand. Is demand at this point relatively responsive or relatively unresponsive to price changes?

2.) For this exercise you will need to first build a graph to these specifications: Draw a downward sloping demand curve with vertical intercept (0,4) and horizontal intercept (8,0). Label the price above which the demand is always elastic. Label the portion of the curve at which a decrease in price will lead to increased revenue. If the price is currently $3, will revenue rise or fall if price drops to $2?
Answer #1

1. Elasticity of demand = % change in quantity/ % change in price

In the mid-point method, the average of the new and old quantity is taken as the base to calculate the % change in quantity and the average of the new and old price is taken as the base to calculate the % change in price.

The quantity goes up from 50 to 100. Change in quantity = 100 - 50 = 50. Base quantity = (100 + 50)/2 = 150/2 = 75

% change in quantity = 50/75 * 100 = 200/3 = 66.67%

The price goes down from 12 to 10. Change in price = 12 - 10 = 2. Base price = (10 + 12)/2 = 22/2 = 11

% change in price = 2/11 * 100 = 200/11 = 18.18%

Elasticity of demand = % change in quantity/ % change in demand = 66.67/18.18 = 3.67

The demad is elastic demand as the value of it is more than 1. Therefore, the demand is relatively responsive to price changes.